# Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model

### v1

@@ Abstract
The diffusion probability model is a state-of-the-art generative model that generates an image by applying a neural network iteratively. Moreover, this generation process is regarded as an algorithm solving a diffusion ordinary differential equation (ODE) or stochastic differential equation (SDE). Based on the analysis of the truncation error of the diffusion ODE and SDE, our study proposes a training-free algorithm that generates high-quality 512 x 512 and 1024 x 1024 images in eight steps, with flexible guidance scales. To the best of my knowledge, our algorithm is the first one that samples a 1024 x 1024 resolution image in 8 steps with an FID performance comparable to that of the latest distillation model, but without additional training. Meanwhile, our algorithm can also generate a 512 x 512 image in 8 steps, and its FID performance is better than the inference result using state-of-the-art ODE solver DPM++ 2m in 20 steps. The result of our algorithm in generating high-quality 512 x 512 images and 1024 x 1024 images with five-step and six-step inference is also comparable to the latest distillation model. Moreover, unlike most distillation algorithms, which achieve state-of-the-art FID performance by fixing the sampling guidance scale, and which sometimes cannot improve their performance by adding inference steps, our algorithm uses a flexible guidance scale on classifier-free guidance sampling. The increase in inference steps enhances its FID performance. Additionally, the algorithm can be considered a plug-in component compatible with most ODE solvers and latent diffusion models. Extensive experiments are performed using the COCO 2014, COCO 2017, and the LAION dataset. Specifically, we validate our eight-step image generation algorithm using the COCO 2014, COCO 2017, and LAION validation datasets with a 5.5 guidance scale and a 7.5 guidance scale, respectively. Furthermore, the FID performance of the image synthesis in 512 x 512 resolution with a 5.5 guidance scale is 15.7, 22.35, and 17.52, meaning it is comparable with the state-of-the-art ODE solver DPM++ in 20 steps, whose best FID performance is 17.3, 23.75, and 17.33, respectively. Further, it also outperforms the state-of-the-art AMED-plugin solver, whose FID performance is 19.07, 25.50, and 18.06. We also apply the algorithm in five-step inference without additional training, for which the best FID performance of our algorithm in COCO 2014, COCO 2017, and LAION is 19.18, 23.24, and 19.61, respectively, which is comparable to the performance of the state-of-the-art AMED Pulgin solver in eight steps, SDXL-turbo in four steps, and the state-of-the-art diffusion distillation model Flash Diffusion in five steps. Then, we validate our algorithm in synthesizing 1024 * 1024 images, whose FID performance in COCO 2014, COCO 2017, and LAION using eight-step inference is 17.84, 24.42, and 19.25, respectively. Thus, it outperforms the SDXL-lightning in eight steps, Flash DiffusionXL in eight steps, and DMD2 in four steps. Moreover, the FID performance of the six-step inference of our algorithm in the 1024 x 1024 image synthesis is 23, which only has a limited distance to the state-of-the-art distillation model mentioned above. We also use information theory to explain the advantage of our algorithm and why it achieves a strong FID performance.
@@ 1 Introduction
The diffusion probability model (DM)[12], a state-of-the-art image generation model with solid mathematics instruction on image sampling, is among the most popular research areas in computer vision. A typical diffusion model, such as stable diffusion (SD)[28], is inspired by the Brownian diffusion process in physics. It consists of two different processes: diffusion and reverse. The diffusion process adds small variants of Gaussian noise to the image step by step to facilitate its gradual conversion into pure Gaussian noise, while the reverse process uses the score matching[37] algorithm to train a neural network studying the gradient of the diffusion process by denoising noisy images[25]. Then, the reverse of the diffusion process occurs. The diffusion process can be described by a diffusion stochastic differential equation (SDE) whose form is similar to the Brownian diffusion SDE. Moreover, the reverse process can be described by an ordinary differential equation (ODE) or an SDE[37]. The ODE is equivalent to the SDE in the reverse process, which generates random variables step by step by transporting the standard Gaussian distribution to a target probability distribution by changing its geometric shape. However, it requires fewer inference steps to sample a high-quality image than solving the SDE, as it prevents involving a random effect during sampling. However, solving diffusion ODEs still requires a large amount of computational resources. Hence, solving this problem remains an active research area. Research related to it can be categorized into two classes. The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process[12, 24, 35, 45, 15, 21]. The second class deems the whole inference process, which uses a neural network iteratively, as a large neural network, and it reduces the inference steps by training a small distillation model[26, 36, 19, 44, 43, 31, 32, 16, 38, 22]. The training-free method typically necessitates a 20-step inference[24], while algorithms requiring additional training can generate an image within four to eight steps. However, our results show that by selecting the proper hyperparameter of the ODE solver and adding a training free diffusion model decorator to exploit the capability of the latent diffusion model, the inference speed can exceed expectations. To that end, we propose a new inference method that can generate 512 x 512 images in eight steps without additional training, and its Frechet Inception Distance (FID)[10] performance surpasses that of the state-of-the-art ODE solver in 20 steps. Furthermore, when applying our algorithm in the five-step inference, it also outperforms the state-of-the-art distillation algorithm in the same number of steps. We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step inference exceeds most of the latest distillation models in eight steps. We perform extensive experiments using the COCO 2014[20], COCO 2017 [20], and LAION datasets[33], and exhibit the generation result in Fig 1. We also use information theory to explain why our algorithm achieves a strong FID result.
@@
Figure 1: The comparison between our algorithm and dpm++ solver in few step inference
@@ 2.1 Diffusion ODE
The image generation process that uses a diffusion model can be regarded as solving a special diffusion SDE or ODE. A typical diffusion model has two processes: forward and reverse. The forward process perturbs the image by adding Gaussian noise to it, which can be described by the following SDE:
@@
where the $w$ is a standard Wiener process. Choosing different $g(t)$ and $f(t)$ corresponds to applying different algorithms. In this study, we use the variant preserve (VP) SDE[37] to describe the diffusion process, which corresponds to the denoised diffusion probability model (DDPM) algorithm[13]. Specifically, its choice is as follows:
$f(t)=-\frac{1}{2}\beta(t),$ (2)
@@
By solving the Eq. 1 step by step with a given $x_{0}$, we obtain a noisy image $x_{t}$. Finally, when applying iteration in $T$ step, we obtain $x_{T}\sim\mathcal{N}(0,\sigma_{T}^{2}I)$. Additionally, using the reparameterization algorithm, we can sample $x_{t}$ within one step[12].
In contrast, the reverse process, which corresponds to generating an image from a pure Gaussian noise, can be described by the following time-dependent ODE[15]:
@@
where $c$ is the condition of input, like the caption that describes the image, $\epsilon_{\theta}(x_{t},t,c)=-\sigma(t)s_{\theta}(x_{t},c)$, where $s_{\theta}(x_{t},c)=\nabla_{x_{t}}\log{p_{\theta}(x_{t},c)}$. Moreover, the score function normally used in VP SDE, it is trained by Eq. 6
@@
and $q_{t}$ is the probability distribution of the random variables sampled by adding noise to the training images in timestep $t$.
Specifically, the $s(t)$ and $\sigma(t)$ is defined by Eq. 7 and Eq. 8:
@@
Using Eq. 5 as a reverse ODE can help exploit the advantage of adapting the DDPM algorithm to the common diffusion framework proposed by Karras [15]. However, the algorithm implementation should also consider whether a neural network can study this input–output pair if the input and output in different timesteps have an obvious gap in the variance when using this framework. So, to let the neural network study properly, we should formulate the equation as
$\epsilon_{\theta}(x_{t},t,c)=c_{out}(\sigma_{t})F_{\theta}(c_{in}(\sigma_{t})x_{t},c_{noise}(\sigma),c),$ (9)
where $F_{\theta}$ is the real function that the neural network studies, $c_{in}$ scales the input to reduce its variance, $c_{out}$ scales the output to use the common framework proposed by Karras, and $c_{noise}(\sigma)$ transforms the variance $\sigma$ to the condition input $t$ of the neural network.
For simplicity, the following section will use $D_{\theta}$ to denote the centrifuge of the possible denoised images predicted by the neural network when given a noisy image $x_{t}$:
$D_{\theta}(x_{t};\sigma)=$ $c_{skip}(\sigma)x_{t}+$ (10)
$c_{out}(\sigma)F_{\theta}(c_{in}(\sigma_{t})x_{t},c_{noise}(\sigma),c),$
where $c_{skip}(\sigma)=1$ in DDPM.
@@ 2.2 Diffusion ODE Solver
Sampling an image from a diffusion model is equivalent to solving the Eq. 5. Moreover, different solvers have been considered in previous research[15, 23, 24]. The inference process approximates the trajectory of the reverse ODE by discretizing its path. The research on reverse ODE solvers attempts to precisely estimate the next step, thereby minimizing the truncation error caused by the discretization. As the culmination of the truncation error will affect the result of synthesis, minimizing the truncation error in each step is equivalent to subtracting the demand for an increased inference step that has caused precise discretization.
Among those solvers, the DPM++ solver[24] is the latest state-of-the-art solver designed specifically for classifier-free guidance (CFG)[11] sampling. It predicts the next step via the following formula:
$x_{t}=\frac{\sigma_{t}}{\sigma_{s}}x_{s}-\sigma_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{\lambda}D_{\theta}(x_{t(\lambda)};\sigma_{t(\lambda)})d\lambda,$ (11)
where $\lambda:=\log{a_{t}}-\log{\sigma_{t}}$ is a log signal-to-noise ratio (SNR) of the noisy image, while the function $(\lambda)$ transforms $\lambda$ to the corresponding $t$.
We use the Taylor formula to obtain an approximation of the Eq. 11, and then there is a first-order DPM++1s solver:
$x_{t}=\frac{\sigma_{t}}{\sigma_{s}}x_{s}-\sigma_{t}e^{h}D_{\theta}(x_{s},\sigma_{s}),$ (12)
in which $h=\lambda_{t}-\lambda_{s}$.
Moreover, by using a multistep method, we obtain the second-order DPM++2m solver:
$x_{t_{i}}=\frac{\sigma_{t_{i}}}{\sigma_{t_{i-1}}}x_{t_{i-1}}-\alpha_{t_{i}}(e^{-h_{i}}-1)D_{i},$ (13)
where $D_{i}$ is as follows:
$D_{i}$ $=(1+\frac{1}{2r_{i}})D_{\theta}(x_{t_{i-1}};\sigma_{t_{i-1}})$ (14)
$-\frac{1}{2r_{i}}D_{\theta}(x_{t_{i-2}};\sigma_{t_{i-2}}),$
and $r_{i}=\frac{h_{i-1}}{h_{i}}$.
@@ 2.3 Diffusion Distillation
The training-free method, which optimizes the truncation error in each step to minimize the demand for precise discretion. Unlike this method, the distillation algorithm[19, 32, 26, 36, 31] considers the whole inference process of applying a neural network iteratively as a large neural network. The distillation algorithm, which normally trains a small neural network, entails reducing its inference steps. However, training a neural network with high-quality content, a fast inference speed, and no additional parameters is difficult. Thus, the distillation algorithm, through the analysis of information theory[17] and experiments, can sometimes reduce the diversity of the generated images, thus reducing the FID performance. This approach may pose a potential risk in the future. AI training requires the use of Internet data, and in some countries, like Japan, all the data can be used in AI training by obeying a series of basic rules. Moreover, when people draw an image via their craft, they will sometimes be affected by the image they have previously watched. If they generate a similar image unconsciously, without directly using a similar work as a reference during creation, such generation will not be deemed plagiarism. However, generative models that use Internet images as training data are always denied by people who claim that if those models generate similar content in their dataset, it should be considered plagiarism. Furthermore, diversity is among the metrics that affect people’s impression of generative content.
@@ 2.4 Latent Diffusion Model
Another method for reducing the calculation resources of the inference is to encode images into a low-dimensional latent space. Stable diffusion[28] performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 PNG image to use the attention mechanism in the U-Net[29] efficiently. One team’s experiment[28] shows that the semantics of the image, which rely on the long-range relationship between different parts of the images, can be modeled by an attention mechanism[40]. In contrast, its perceptual information, which makes the image appear crisp, is strongly linked to local regions and can be modeled by a convolutional neural network (CNN)[18]. Hence, the stable diffusion model first encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder[2] to let the following inference ignore the perceptual detail. Then, it performs diffusion inference in latent space via a U-Net[29] with an attention mechanism to model the semantic information. Experiments show that compared with vector quantify VAE[41], the normal $\beta$-VAE with a small KL-term constraint generates higher-quality images.

### v2

@@ Abstract
The diffusion model is a state-of-the-art generative model that samples images by applying a neural network iteratively. However, the original sampling algorithm requires substantial computation cost, and reducing the sampling step is a prevailing research area. To cope with this problem, one mainstream approach is to treat the sampling process as an algorithm that solves an ordinary differential equation (ODE). Our study proposes a training-free inference plugin compatible with most few-step ODE solvers. To the best of my knowledge, our algorithm is the first training-free algorithm to sample a 1024 x 1024-resolution image in 6 steps and a 512 x 512-resolution image in 5 steps, with an FID result that outperforms the SOTA distillation models and the 20-step DPM++ 2m solver, respectively. Based on analyses of the latent diffusion model’s structure, the diffusion ODE, and the Free-U mechanism, we explain why specific hyperparameter couplings improve stability and inference speed without retraining. Meanwhile, experimental results also reveal a new design space of the latent diffusion ODE solver. Additionally, we also analyze the difference between the original diffusion model and the diffusion distillation model via an information-theoretic study, which shows the reason why the few-step ODE solver designed for the diffusion model can outperform the training-based diffusion distillation algorithm in few-step inference. The tentative results of the experiment prove the mathematical analysis.
@@ 1 Introduction
The diffusion probability model (DM) proposed by Ho et al. (2020a), a state-of-the-art image generation model with solid mathematics instruction on image sampling, is among the most popular research areas in computer vision. A typical diffusion model, such as stable diffusion (SD) (Rombach et al. (2022)), is inspired by the Brownian diffusion process in physics. It consists of two different processes: diffusion and reverse. The diffusion process adds small variants of Gaussian noise to the image step by step to convert an image into a pure Gaussian noise, while the reverse process uses the score matching algorithm (Song et al. (2020b)) to train a neural network studying the gradient of the diffusion process by denoising noisy images (Luo (2022)).
Like the Brownian process in physics, the diffusion process can be described by a diffusion stochastic differential equation (SDE). Moreover, the research proposed by Song et al. (2020b) shows that there exists an ordinary differential equation (ODE) to describe the reverse process. And using the existing mathematical tool for ODE solving enhances the capability of the diffusion model to sample an image.
However, compared to typical algorithms such as generative adversarial networks and variational autoencoders, solving diffusion ODEs still requires substantial computational resources. Hence, coping with this problem remains an active research area.
Research related to it can be categorized into two classes. The first class proposes a training-free algorithm aimed at providing a good SDE and ODE solver tailored to the diffusion process (Ho et al. (2020a); Lu et al. (2025); Song et al. (2020a); Zhao et al. (2023); Karras et al. (2022); Liu et al. (2022)). The second class deems the whole inference process, which uses a neural network iteratively, as a large neural network, and the algorithm reducing the inference steps is equivalent to train a small distillation model (Luo et al. ; Song & Dhariwal (2023); Lin et al. (2024); Yin et al. (2024b; a); Salimans & Ho (2022); Sauer et al. (2024); Kim et al. (2023); Song et al. (2023); Liu et al. (2023)). The training-free method typically necessitates a 20-step inference, like the DPM++ solver proposed by Lu et al. (2025), while algorithms requiring additional training can generate an image within four to eight steps. However, our results show that by selecting the appropriate hyperparameter for the ODE solver and adding a training-free diffusion model decorator to leverage the latent diffusion model’s capabilities, the inference speed can exceed expectations.
To that end, we propose a new inference method that generates 512 x 512 images in 5 steps without additional training, and its Frechet Inception Distance (FID) (Heusel et al. (2017)) performance surpasses both the DPM++ 2m solver and the SOTA distillation model in 20 steps and 8 steps, respectively. We also apply our algorithm to 1024 x 1024 image generation, and the FID performance of the eight-step and six-step inference exceeds most of the latest distillation models. We perform extensive experiments using the COCO 2014 (Lin et al. (2014)), COCO 2017 (Lin et al. (2014)), and LAION datasets (Schuhmann et al. (2022)), and exhibit the generation result in Fig 1. We also use information theory to explain why our algorithm achieves a strong FID result.
@@
Figure 1: The comparison between our algorithm and UniPC solver in few step inference
@@ 2.1 Diffusion ODE
The image-generation process of the diffusion model can be deemed as solving a special diffusion ODE. A typical diffusion model has two processes: forward and reverse. The forward process perturbs the image by adding Gaussian noise to it, which can be described by the following SDE:
@@
where the $w$ is a standard Wiener process. Choosing different $g(t)$ and $f(t)$ corresponds to applying different algorithms. In this study, we use the variant preserve (VP) SDE proposed by Song et al. (2020b) to describe the diffusion process, which corresponds to the denoised diffusion probability model (DDPM) algorithm (Ho et al. (2020b)). Specifically, its choice is as follows:
$f(t)=-\frac{1}{2}\beta(t)=\frac{d\log{\alpha_{t}}}{dt},$ (2)
@@
Explicitly, in the diffusion process mentioned above, the variant of noise added at step $t$ is $\beta_{t}$. Explicilty, $\beta_{max}$ and $\beta_{min}$ can be determined arbitrarily if it can ensure $p(x_{T}|x_{0})$ is a standard Gaussian distribution. In the timestep $t$, the noisy feature $x_{t}$ with given $x_{0}$ belongs to a Gaussian distribution described in the eq.˜5.
$p(x_{t}|x_{0})=\mathcal{N}(\alpha(t)x_{0};\sigma^{2}(t)I).$ (5)
Typically, to transform an image into a standard Gaussian noise, the DDPM algorithm requires $T=1000$ steps, and $\alpha_{T}$ is closed to the zero. Additionally, using the reparameterization algorithm, we can sample $x_{t}$ in a single step (Ho et al. (2020a)).
In contrast, the reverse process, which corresponds to generating an image from a pure Gaussian noise, can be described by the following time-dependent ODE proposed by Karras et al. (2022):
@@
where $c$ is the condition of input, like the caption that describes the image, and the $\epsilon_{\theta}$ is modeled by a neural network. The $\epsilon_{\theta}$ has a relation between the score function, which is described in the eq.˜7:
$\epsilon_{\theta}(x_{t},t,c)=-\sigma(t)score_{\theta}(x_{t},c),$
$score_{\theta}(x_{t},c)=\nabla_{x_{t}}\log{p_{\theta}(x_{t},c)},$ (7)
And the $p_{\theta}(x_{t},c)$ is the probability distribution of noisy images $x_{t}$ modeled implicitly by the neural network with parameters $\theta$. Moreover, the score function normally used in VP SDE is trained by eq.˜8
@@
and $q_{t}(x)$ is the probability distribution of the random variables sampled by adding noise to the training images in timestep $t$, and $q_{0}$ means the distribution of clean images.
In the eq.˜6, the $s(t)$ and $\sigma(t)$ is defined by eq.˜9 and eq.˜10:
@@
Deeming the loss of the DDPM algorithm as a score function described by the eq.˜8 can help exploit the advantage of adapting the original diffusion model to the common diffusion ODE framework proposed by Karras et al. (2022).
@@ 2.2 Diffusion ODE Solver
Sampling an image from a diffusion model is equivalent to solving the eq.˜6. Via the research from Lu et al. (2022), when given a noisy input $x_{s}$ and a target timestep $t$, the correct noisy feature output $x_{t}$ can be gained by solving the eq.˜11 :
$x_{t}=\frac{\alpha({t})}{\alpha(s)}x_{s}-\alpha_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{-\lambda}\epsilon_{\theta}(x_{t(\lambda)},\lambda)d\lambda,$ (11)
where $\lambda_{t}$ is the log Signal-to-Noise Ratio (SNR) of the noisy feature $x_{t}$.
Sometimes, to cope with the high guidance scale on the classifier free guidance algorithm (CFG) (Ho & Salimans (2022)), researchers tend to replace eq.˜11 with the eq.˜12 to get a better result (Lu et al. (2025)):
$x_{t}=\frac{\sigma({t})}{\sigma(s)}x_{s}-\sigma_{t}\int^{\lambda_{t}}_{\lambda_{s}}e^{\lambda}\hat{x}_{\theta}(x_{t(\lambda)},\lambda)d\lambda,$ (12)
where $\hat{x}_{\theta}$ is a clean feature predicted indirectly by the neural network.
By using the Taylor Expand in the $\epsilon_{\theta}$ or $x_{\theta}$ (Lu et al. (2022; 2025)), we get the following eq.˜13:
$x_{t}$ $=\frac{\alpha({t})}{\alpha(s)}x_{s}-\alpha_{t}\Sigma_{n=0}^{k-1}{\epsilon}^{(n)}_{\theta}(x_{t(\lambda)},\lambda)\int^{\lambda_{t}}_{\lambda_{s}}e^{-\lambda}\frac{(\lambda-\lambda_{s})^{n}}{n!}d\lambda$
$+\mathcal{O}(h_{i}^{k+1}),$ (13)
However, calculating the neural networks’ deviation requires substantial computational resources. Researchers typically use the Runge-Kutta method (Lu et al. (2025); Zhao et al. (2023)) to approximate the deviation.
Explicitly, the DPM++ 2s algorithm approximate $x_{t}$ via the following formula:
$x_{t}$ $=\frac{\sigma({t})}{\sigma(s)}x_{s}-\alpha_{t}(e^{-h_{t}}-1)\hat{x}_{\theta}(x_{s},s)$
$+\frac{1}{2r_{t}}*\alpha_{t}(e^{-h_{t}}-1)*(\hat{x}_{\theta}(\mu,m)-\hat{x}_{\theta}(x_{s},s)),$ (14)
in which $r_{t}$ can be selected arbitarily between 0 and 1, and $h_{i}$, $m$ and $\mu_{i}$ is given below:
$h_{t}$ $=\lambda_{t}-\lambda_{s},$ (15)
$r_{t}$ $=\frac{\lambda_{m}-\lambda_{s}}{h_{i}},$ (16)
$\mu_{t}$ $=\frac{\sigma_{m}}{\sigma_{s}}x_{s}-\alpha_{m}(e^{-r_{t}h_{t}}-1)\hat{x}_{\theta}(x_{s},s),$ (17)
To get the first-order deviation of the $\epsilon_{\theta}$, the number of function evaluations (NFE) in DPM++ 2s is 2. While DPM++ 2m utilizes the latest cached output in the previous inference to calculate the approximation of the first-order deviation, and hence only requires 1 NFE, which is a typical multistep algorithm in ODE solving. Simultaneously, the UniPC and DPM V3 solver cached all of the previous output to correct and predict a high-order deviation.
The eq.˜6 describes a continuous curved trajectory, and the ODE solver approximates it by discretizing the path with a series of lines. Research on reverse ODE solvers aims to precisely estimate the next step, thereby minimizing truncation error from pool discretization in the few-step inference. On the other hand, the discrete method proposed by Karras et al. (2022) does not aim at design a new ODE solving algorithm, but proposes a choice of moving distance $h_{t}$ to enhance the sample quality. Via the analysis of the Karras’ research, the $h_{t}$ should be large at the beginning of the inference, and then decrease monotonically.
@@ 2.3 Diffusion Distillation
The training-free method optimizes the truncation error at each step and thus minimizes the need for additional inference steps for the precise discretization. Unlike this method, the distillation algorithm (Lin et al. (2024); Sauer et al. (2024); Luo et al. ; Song & Dhariwal (2023); Salimans & Ho (2022)) considers the whole inference process of applying a neural network iteratively as an extensive neural network. The distillation algorithm, which typically trains a small neural network, reduces its inference steps. However, training a neural network with high-quality content, fast inference, and no additional parameters is difficult. Thus, the distillation algorithm, through the analysis of information theory (Li & Zhang (2025)) and experiments, can sometimes reduce the diversity of the generated images, and affect the FID performance. This approach may pose a potential risk in the future. AI training requires the use of Internet data, and in some countries, like Japan, all the data can be used in AI training by obeying a series of basic rules. However, when a human draws an image via their craft, they will sometimes be affected by the image they have previously seen. If they generate a similar image unconsciously, without directly using a similar work as a reference during creation, such a generation will not be deemed plagiarism. Compared to this, generative models trained on Internet images are often criticized, with people claiming that if those models generate content similar to their dataset, it should be considered plagiarism. In such a situation, diversity is a metrics that affect people’s impression of generative content.
@@ 2.4 Latent Diffusion Model
Another method for reducing the computational requirements of inference is to encode images into a low-dimensional latent space. Stable diffusion performs inference in a 64 x 64 x 4 latent space to generate a 512 x 512 x 4 PNG image. One experiment from Rombach et al. (2022) shows that the semantics of the image, which rely on the long-range relationship between different parts of the images, can be modeled by an attention mechanism (Vaswani et al. (2017)). In contrast, its perceptual information, which provides the image with sufficient real details, is strongly tied to local regions and can be modeled by a convolutional neural network (CNN) (e.g. Li et al. (2021)). Hence, the stable diffusion model first encodes the image into a low-dimensional latent space via a $\beta$-variational autoencoder (Burgess et al. (2018)) to let the following inference ignore the perceptual detail. Then, it performs diffusion inference in latent space via a U-Net (Ronneberger et al. (2015)) with an attention mechanism to model the semantic information. Experiments show that compared with vector quantify VAE (Wu & Flierl (2020)), the normal $\beta$-VAE with a small KL-term constraint generates higher-quality images.